OpenStudy (anonymous):

John took money out of his account. He bought a radio for $108. He used 5/9 of what was left for presents and then put 5/6 of the leftover into another account and gave away the remaining $80 to charity. How much did he start with?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

found you

OpenStudy (anonymous):

Ok, let's denote $ he start with as X

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Since he bought the radio he has less money. $108 less money so x-108 make sense so far?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let's go to next part of the question. " He used 5/9 of what was left for presents" The leftover is x-108. So 5/9 of left over is 5/9(x-108 ) good?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

So money left after buying the radio is( x-180) and he spend 5/9 of it. So we subtract. \[(x-180)-\frac{5}{9}(x-180)\] We can factor to make it look neat \[(x-180)(1-\frac{5}{9})\] \[(x-180)(\frac{4}{9})\] So this is how much money he has left after he bought the radio,use money for the present. Does this make sense?

OpenStudy (anonymous):

no--back up, to the (x-108) -5/9(x-108)

OpenStudy (anonymous):

(x-108) is how much he has left after buying the radio. He used 5/9 of left over to buy present. So we do (x-108) -5/9(x-108) to get new leftover

OpenStudy (anonymous):

Would that be -5/9x + 60

OpenStudy (anonymous):

I see you multiplied it ,yes

OpenStudy (anonymous):

5/9 * 108

OpenStudy (anonymous):

that's correct so subtract -5/9x + 60 from (x-108) to get new left over x-108-5/9x+60

OpenStudy (anonymous):

I factored it, instead of multiplying out

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

what next

OpenStudy (anonymous):

"then put 5/6 of the leftover into another account" \[ (x-108-\frac{5}{9}x+60) -\frac{5}{6}(x-108-\frac{5}{9}x+60)\]

OpenStudy (anonymous):

Any question?

OpenStudy (anonymous):

whereis the 80 that is left

OpenStudy (anonymous):

I am not there yet

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

But so far so good right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

" gave away the remaining $80 to charity" This part tell us that he has 80 remaining after buying radio,buying present, putting money in another account. So we may set the previous equation equals to 80 \[(x−108−\frac{5}{9}x+60)−\frac{5}{6}(x−108−\frac{5}{9}x+60)=80\] Now you are all set to solve for x, just some algebra

OpenStudy (anonymous):

don't lave me now__I have stressed over this thing enough

OpenStudy (anonymous):

You solve for x; I will too then we will compare result

OpenStudy (anonymous):

* Tip First add like term inside the parenthesis before distributing

OpenStudy (anonymous):

It isn't working!!

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

it is supposed to come out to 1188 and I'm not getting it.

OpenStudy (anonymous):

That's what I got

OpenStudy (anonymous):

did you do the sectios separate or together

OpenStudy (anonymous):

I do not have a math gene in my body

OpenStudy (anonymous):

\[\left(x-108-\frac{5}{9}x+60\right)-\frac{5}{6}\left(x-108-\frac{5}{9}x+60\right)\text{==}80\] Add like term inside the parethesis \[(-48+\frac{4 x}{9})-\frac{5}{6}(-48+\frac{4 x}{9})==80\] \[(-48+\frac{4 x}{9}+40-\frac{10 x}{27})==80\] \[-8+\frac{2 x}{27}=80\] \[\frac{2 x}{27}=88\] \[{2 x}=2376\] \[{2 x}=2376\] X=1128

OpenStudy (anonymous):

2x=2376 x=1188

OpenStudy (anonymous):

This I understand___I just had to see my error__it was the algebra--I did not combine the x's properly.

OpenStudy (anonymous):

You have lifted a weight off my shoulders.

OpenStudy (anonymous):

Yeah, be careful with algebra, use calculator if you are allowed

OpenStudy (anonymous):

I will--exam is tomorrow

OpenStudy (anonymous):

What grade?

OpenStudy (anonymous):

It is a college course for elementary education. I need it to graduate next month.

OpenStudy (anonymous):

It is online and I don't get much feedback.

OpenStudy (anonymous):

Good luck on the test tomorrow.

OpenStudy (anonymous):

Thank you again. good nite